OpenStudy (anonymous):

Assume that a procedure yields a binomial distribution with a trial repeated n times. Use the binomial probability formula to find the probability of x successes given the probability p of success on a single trial. Round to three decimal places. n = 4, x = 3, p = 1/6

OpenStudy (dumbcow):

=4c3 (1/6)^3*(5/6)

OpenStudy (anonymous):

\[P(x=3)=\dbinom{4}{3}(\frac{1}{6})^3(\frac{5}{6})^1\]

OpenStudy (anonymous):

\[=4\times \frac{5}{6^4}=\frac{20}{1296}=\frac{5}{324}\]\]